Hi.
I'm checking to see if I solved this problem correctly.
Solve each system by by method.
{2/3x-1/4y=-8
1/2x-3/8y=-9
thanks
H.C.

Answer by eperette(173)

I will solve using elimination method
2/3x-1/4y=-8
1/2x-3/8y=-9
I always find it easier to get rid of fractions first...you can do so by
(1) finding common denominator of fractions in each equation (multiply denominators times each other of each equation)
12 ( 2/3 x - 1/4 y = -8)
16 ( 1/2 x - 3/8 y = -9)
****************************
24/3 x - 12/4 y = 12(-8)
16/2 x - 48/8 y = 16(-9)
****************************
8x - 3y = -96
8x - 6y = -144
****************************
Now subtract second equation from first
(8x - 3y = -96
-(8x - 6y = -144)
0x + 3y = 48
3y/3 = 48/3
y= 16
**************************
replace y into one equation to find x
8x - 3y = -96
8x - 3(16) = -96
8x - 48 = -96
8x - 48+48 = -96 + 48
8x - -48
8x/8 = -48/8
x= -6
*************************
so answer is (-6, 16)
*********************************************
LETS CHECK:
2/3x-1/4y=-8
2/3(-6) - 1/4(16)
-12/3 - 16/4
-4 - 4
-8
1/2x-3/8y=-9
1/2(-6) - 3/8(16)
-6/2 - 48/8
-3 - 6
-9
the coordinate point works for both equations so it is right